# What exactly does Bekenstein bound?

### v2

@@ 1.1 Our contributions
What if one also restrains the encoder’s spatial domain and energy budget? Does it help constrain the zero-bit capacity? In previous studies of the Bekenstein bound, such subtleties did not arise because people sought a the Bekenstein bound that works for any state. In other words, it was implicitly assumed that the encoding is not restricted at all. In the communication setting, however, it is physically reasonable to consider physical restrictions on the encoder. Then it is sensible to expect that the constraints on the encoder and the decoder should both be manifest in the Bekenstein bound.
@@ 2 Motivation: The Bekenstein bound on Accessible Information
Therefore, it is the same as Casini’s bound (1.2) up to the regularization. While the vacuum-subtracted energy may be natural, vacuum-subtracted entropy is more ad hoc. The regularized quantities in (2.5), in contrast, are always well-defined as relative entropies. While Casini’s entropy bound $\langle K_{B}\rangle_{\rho}$ is linear in the quantum state $\rho$, (2.5) is linear in $p_{X}$, which describes different sources of classical information. As compared to Casini’s bound on entropy, the Holevo information is operationally meaningful and nonnegative. Bousso further considered signals that are classical enough that the codewords are more entropic than the vacuum,
@@ 3.1 Communication with the Unruh channel
We have identified the Hilbert space of the right Rindler wedge as Bob’s Hilbert space $\mathcal{H}_{B}$.999 Note that the quantum field theory Hilbert space $\mathcal{H}$ is generally nonfactorizable. However, when restricted to the subspace of particles in a single mode, the restricted Hilbert space $\mathcal{H}$ does factorize into $\mathcal{H}=\mathcal{H}_{L}\otimes\mathcal{H}_{R}$.
@@
Before proceeding, let us briefly comment on the physicality of this communication channel. Note that we did not model the time evolution of the mode or give Bob any time window for decoding. Therefore we are describing a simplified and idealized communication scenario, where there Bob has infinite computational power to decode the message in a short amount of time. This is accepted so long as we are testing the validity of Bekenstein bound from the purely information-theoretic perspective.
Unruh modes are often used to approximate Minkowski modes in the so-called single-mode approximation alsing2003teleportation , which is sometimes problematic bruschi2010unruh . We assume Alice can prepare the Unruh modes directly rather than using them as an approximation. They are convenient because their Bogoliubov transformation to the Rindler modes is easier to analyze than the one for Minkowski modes. However, the Unruh modes are less regular than the Minkowski modes. For an inertial observer, they are non-monochromatic, ill-localized, and rapidly oscillatoring near the Rindler horizon. Hence, they are much more difficult to prepare physically.
@@ 3.2 A lightning review of channel capacities
$\chi(\mathcal{M}):=\max_{\{p_{x},\psi_{x}\}}\left[S\left(\sum_{a}p_{x}\mathcal{M}(\psi_{x})\right)-\sum_{x}p_{x}S(\mathcal{M}(\psi_{x}))\right]\ ,$ (3.11)
@@
where $\psi\in\mathcal{H}_{A}\otimes\mathcal{H}_{A^{\prime}}$, $\mathcal{M}(\psi)\in\mathcal{H}_{A^{\prime}}\otimes\mathcal{H}_{B}$.
@@
We are also interested in the classical and quantum capacities of a channel when Alice and Bob are given free entanglement as resource. In that setting, Alice encodes her messages into her half of a set of shared Bell pairs and sends the encoded halves to Bob through the channel. Bob will try to decode Alice’s messages with the help of his half of the Bell pairs. Since entanglement can only help them achieve better communication rates, the entanglement-assisted capacities are generally higher than without the assistance. The capacities are defined in a similar way, so we shall not repeat the details. We will use two useful facts here. Bennett et al bennett1999entanglement ; bennett2002entanglement showed that the entanglement-assisted classical capacity is always given by a single-letter formula that doesn’t involve multiple copies,
@@
Note that $C_{E}(\mathcal{M})$ also equals to
@@
  • Covariance. A covariant channel has its input and output transform covariantly under the same unitary group. More precisely, we say that a channel $\mathcal{M}:\mathcal{P}(\mathcal{H}_{A})\to\mathcal{P}(\mathcal{H}_{B})$ is $G$-covariant with respect to representations $r_{1}:G\to\mathrm{GL}(\mathcal{H}_{A}),r_{2}:G\to\mathrm{GL}(\mathcal{H}_{B})$, if $\mathcal{M}(r_{1}(g)\rho\,r_{1}(g)^{\dagger})=r_{2}(g)\mathcal{E}(\rho)r_{2}(g)^{\dagger},\ \forall g\in G,\rho\in\mathcal{P}(\mathcal{H}_{A})$. The claim is that the Unruh channel is $\mathrm{SU}(d)$-covariant with respect to the fundamental representation $U(g)$ in the input space and the isometrically mapped representation $VU(g)V^{\dagger}$ in the output space.
@@
    The upshot is that although the quantum capacity is generally superadditive, but it is additive for degradable channels devetak2005capacity . We therefore have a single-letter formula,
@@
  • Hadamard. The Unruh channel is a Hadamard channel bradler2011infinite , [^12] The channel action in the Choi representation takes the form of a Hadamard product. meaning that its complement channel is entanglement-breaking. [^13] Entanglement-breaking channels output separable states for any inputs entangled with a reference horodecki2003entanglement . They have additive classical capacities and zero quantum capacity. Here $\mathcal{N}^{c}$ being entanglement-breaking follows from the fact that its transpose $\mathcal{T}\circ\mathcal{N}^{c}$ is also a quantum channel. Hadamard channel is a special instance of a degradable channel: [^14] Hence, in fact, the Unruh channel is both degradable and transpose degradable. Its degrading map can be decomposed as two maps, where the first one maps the channel output to a classical description and the second one prepares the complementary output according to the classical description.
@@ 4 No Bekenstein bound for zero-bits
One caveat that might undermine our result is that this violation is an artifact of choosing the Unruh mode of various species as codewords. We make this choice for the convenience of additivity. Since the Unruh mode is singular at the horizon, it is technically possible that Unruh mode could be fine-tuned for this zero-bit species problem. To check the robustness of our example, one needs to adopt and perhaps generalize the continuity statements concerning approximate additive channels sutter2017approximate ; leditzky2018approaches , which claim that channels close in diamond norm to an additive one have capacities close to the additive counterparts.
@@ 5 Restraining the encoder
As compared to the more well-known bound of Casini (1.2), this version of the Bekenstein bound has the upshot that both sides are non-negative, and the bound has manifestly the canonical form (1.1). However, the operational meaning of this conditional mutual information has not been elucidated by Blanco-Casini. [^24] One possible interpretation is that the (classical or quantum) communication resource that needed for the task of redistributing the sysmte $A\cap B$ from $A$ to $B$ devetak2008exact .
Let us consider a communication scenario in which Alice encodes her messages by acting locally in $A$. This is illustrated in Fig. 4(b). Considering the time evolution, we let Alice encoder her state in some half space in the past and demand that the future light cone intersects Bob’s $t=0$ slice at the subregion $A\cap B$. Therefore, Alice’s encoding region and Bob’s decoding region overlap over a strip region $A\cap B$ at $t=0$. Compared to the Unruh channel, we have now constrained Alice’s encoding operations. The receiver Bob still sees the right Rindler wedge defined by the half-space $B$.
@@
$\max_{\psi\in\mathcal{H}_{A}\otimes\mathcal{H}_{R}}I(B:R)_{\mathcal{M}(\psi)}\leq 2\pi L\langle\mathcal{H}\rangle_{\mathcal{M}(\psi^{*})}=:2\pi LE\ ,$ (5.4)
@@
Since we have considered the case of restraining only the decoder and the case of restraining both the encoder and the decoder, one could wonder what if the Bekenstein bound applies if only the encoder is restrained. It turns out that it is still crucial to restrain the decoder for the Bekenstein bound to hold. We revisit an argument due to Page in Appendix B, which essentially shows that no Bekenstein bound could work otherwise.
@@ Appendix A No Bekenstein bound for $\alpha$-bits
Hayden-Penington proposed a generalization of a zero-bit, called an $\alpha$-bit ($0\leq\alpha\leq 1$hayden2020approximate . An $\alpha$-bit interpolates between a zero-bit ($\alpha=0$) and a qubit ($\alpha=1$). The intuition is that while a noisy channel might only be able to transmit a decent amount of noisy qubits, we can still make good use of them by distilling an $\alpha$-fraction of each qubit. Just like the zero-bits, the so-obtained $\alpha$-bits can be used as tighter communications resources in various tasks. [^27] For instance, the $\alpha$-bit dense-coding resource inequality reads, $1\;\alpha\mathrm{-bit}+1\;\mathrm{ebit}\geq(1+\alpha)\;\mathrm{cbits}$.
Let us define $\alpha$-bits in the communication scenario. (cf. Section 3 of hayden2020approximate for the precise definition.) Recall that the zero-bit capacity $Q_{0}(\mathcal{M})$ characterizes how much information can the channel transmit if we only require Bob to decode a two-dimensional (or any constant) subspace. The $\alpha$-bit capacity $Q_{\alpha}(\mathcal{M})$ is defined similarly by requiring Bob to decode, with sufficiently small error at large enough $n$, a $k^{n\alpha}$-dimensional subspace of the $k^{n}$-dimensional input message transmitted by $\mathcal{M}^{\otimes n}$. We then say that a channel $\mathcal{M}$ transmits a $Q_{\alpha}(\mathcal{M})$ amount of $\alpha$-bits (asymptotically).
@@
It is natural to ask for which $\alpha$ values, there is a violation of the Bekenstein bound $\langle K_{B}\rangle=\beta E$. Given a fixed $\beta\omega$, we find that there is a violation for
@@
As a sanity check, we note that there’s no lower bound for the range of Bekenstein-violating $\alpha$ values, so the zero-bits always violate the bound at some temperature. On the other hand, the upper bounds are strictly less than $1$ because $Q(\mathcal{N})<\beta E$, meaning that qubits obey the Bekenstein bound. Also, the maximal $\alpha$ that satisfies (A.3) can be arbitrarily close to $1$ at a sufficiently large $d$, because in this limit $Q(\mathcal{N})$ tends to $\beta E$ from below. In Fig. 5, we plotted a “phase diagram” that illustrates (A.3) for $d=2^{6}$.
@@ Appendix D Zero-bit dense coding
$X:=\sum_{0\leq j<d}\left|j+1\mod\ d\right\rangle\!\!\left\langle j\right|,\quad Z:=\sum_{0\leq j<d}e^{i2\pi j/d}\left|j\right\rangle\!\!\left\langle j\right|$ (D.1)

### v3

@@ 1.1 Our contributions
What if one also restrains the encoder’s spatial domain and energy budget? Does it help constrain the zero-bit capacity? In previous studies of the Bekenstein bound, such subtleties did not arise because people sought the Bekenstein bound that works for any state. In other words, it was implicitly assumed that the encoding is not restricted at all. In the communication setting, however, it is physically reasonable to consider physical restrictions on the encoder. Then it is sensible to expect that the constraints on the encoder and the decoder should both be manifest in the Bekenstein bound.
@@ 2 Motivation: The Bekenstein bound on Accessible Information
Therefore, it is the same as Casini’s bound (1.2) up to the regularization. While the vacuum-subtracted energy may be natural, vacuum-subtracted entropy is more ad hoc. The regularized quantities in (2.5), in contrast, are always well-defined as relative entropies. While Casini’s entropy bound $\langle K_{B}\rangle_{\rho}$ is linear in the quantum state $\rho$, (2.5) is linear in $p_{X}$, which describes different sources of classical information. As compared to Casini’s bound on entropy, the Holevo information is operationally meaningful and nonnegative. Bousso further considered signals that are classical enough that the codewords are more entropic than the vacuum, [^7] In any relativistic quantum field theory, the vacuum is infinitely entangled, so it has a UV-divergent entanglement entropy. Here the inequality must be understood as after subtracting the same UV-divergent piece from both sides.
@@ 3.1 Communication with the Unruh channel
We have identified the Hilbert space of the right Rindler wedge as Bob’s Hilbert space $\mathcal{H}_{B}$.101010 Note that the quantum field theory Hilbert space $\mathcal{H}$ is generally nonfactorizable. However, when restricted to the subspace of particles in a single mode, the restricted Hilbert space $\mathcal{H}$ does factorize into $\mathcal{H}=\mathcal{H}_{L}\otimes\mathcal{H}_{R}$.
@@
Before proceeding, let us briefly comment on the physicality of this communication channel. Note that we did not model the time evolution of the mode or give Bob any time window for decoding. Therefore we are describing a simplified and idealized communication scenario, where there Bob has infinite computational power to decode the message in a short amount of time. This is acceptable so long as we are testing the validity of the Bekenstein bound from the purely information-theoretic perspective.
Unruh modes are often used to approximate Minkowski modes in the so-called single-mode approximation alsing2003teleportation , which is sometimes problematic bruschi2010unruh . We assume Alice can prepare the Unruh modes directly rather than using them as an approximation. Also, we restrict Alice to encode the message with single-particle excitations. These assumptions are convenient because their Bogoliubov transformation to the Rindler modes is easier to analyze than the one for Minkowski modes. This technical assumption helps to facilitate channel capacity calculations. However, we should mention other the Unruh modes are less regular than the Minkowski modes. For an inertial observer, they are non-monochromatic, ill-localized, and rapidly oscillatoring near the Rindler horizon. Hence, they are much more difficult to prepare physically.
@@ 3.2 A lightning review of channel capacities
$\chi(\mathcal{M}):=\max_{\{p_{x},\psi_{x}\}}\left[S\left(\sum_{x}p_{x}\mathcal{M}(\psi_{x})\right)-\sum_{x}p_{x}S(\mathcal{M}(\psi_{x}))\right]\ ,$ (3.11)
@@
where $\psi\in\mathcal{H}_{A}\otimes\mathcal{H}_{A^{\prime}}$, $\mathcal{M}(\psi)\in\mathcal{H}_{A^{\prime}}\otimes\mathcal{H}_{B}$, and $A^{\prime}$ is an auxiliary system isomorphic to $A$ that acts as a reference system to help capture the quantum information in $A$.
@@
We are also interested in the classical and quantum capacities of a channel when Alice and Bob are given free entanglement as a resource. In that setting, Alice encodes her messages into her half of a set of shared Bell pairs and sends the encoded halves to Bob through the channel. Bob will try to decode Alice’s messages with the help of his half of the Bell pairs. Since entanglement can only help them achieve better communication rates, the entanglement-assisted capacities are generally higher than without the assistance. The capacities are defined in a similar way, so we shall not repeat the details. We will use two useful facts here. Bennett et al. bennett1999entanglement ; bennett2002entanglement showed that the entanglement-assisted classical capacity is always given by a single-letter formula that doesn’t involve multiple copies,
@@
We can also rewrite the formula as
@@
  • Covariance. A covariant channel has its input and output transform covariantly under the same unitary group. More precisely, we say that a channel $\mathcal{M}:\mathcal{P}(\mathcal{H}_{A})\to\mathcal{P}(\mathcal{H}_{B})$ is $G$-covariant with respect to representations $r_{1}:G\to\mathrm{GL}(\mathcal{H}_{A}),r_{2}:G\to\mathrm{GL}(\mathcal{H}_{B})$, if $\mathcal{M}(r_{1}(g)\rho\,r_{1}(g)^{\dagger})=r_{2}(g)\mathcal{M}(\rho)r_{2}(g)^{\dagger},\ \forall g\in G,\rho\in\mathcal{P}(\mathcal{H}_{A})$. The claim is that the Unruh channel is $\mathrm{SU}(d)$-covariant with respect to the fundamental representation $U(g)$ in the input space and the isometrically mapped representation $VU(g)V^{\dagger}$ in the output space.
@@
    The upshot is that although the quantum capacity is generally superadditive, it is additive for degradable channels devetak2005capacity . We therefore have a single-letter formula,
@@
  • Hadamard. The Unruh channel is a Hadamard channel bradler2011infinite , [^13] The channel action in the Choi representation takes the form of a Hadamard product. meaning that its complement channel is entanglement-breaking. [^14] Entanglement-breaking channels output separable states for any inputs entangled with a reference horodecki2003entanglement . They have additive classical capacities and zero quantum capacity. Here $\mathcal{N}^{c}$ being entanglement-breaking follows from the fact that its transpose $\mathcal{T}\circ\mathcal{N}^{c}$ is also a quantum channel. A Hadamard channel is a special instance of a degradable channel: [^15] Hence, in fact, the Unruh channel is both degradable and transpose degradable. Its degrading map can be decomposed as two maps, where the first one maps the channel output to a classical description and the second one prepares the complementary output according to the classical description.
@@ 4 No Bekenstein bound for zero-bits
One caveat that might undermine our result is that this violation is an artifact of choosing the Unruh mode of various species as codewords. We make this choice for the convenience of additivity. Since the Unruh mode is singular at the horizon, it is technically possible that the Unruh mode could be fine-tuned for this zero-bit species problem. To check the robustness of our example, one needs to adopt and perhaps generalize the continuity statements concerning approximate additive channels sutter2017approximate ; leditzky2018approaches , which claim that channels close in diamond norm to an additive one have capacities close to the additive counterparts.
@@ 5 Restraining the encoder
As compared to the more well-known bound of Casini (1.2), this version of the Bekenstein bound has the upshot that both sides are non-negative, and the bound has manifestly the canonical form (1.1). However, the operational meaning of this conditional mutual information has not been elucidated by Blanco-Casini. [^25] One possible interpretation is that the (classical or quantum) communication resource that needed for the task of redistributing the system $A\cap B$ from $A$ to $B$ devetak2008exact .
Let us consider a communication scenario in which Alice encodes her messages by acting locally in $A$. This is illustrated in Fig. 4(b). Considering the time evolution, we let Alice encode her state in some half space in the past and demand that the future light cone intersects Bob’s $t=0$ slice at the subregion $A\cap B$. Therefore, Alice’s encoding region and Bob’s decoding region overlap over a strip region $A\cap B$ at $t=0$. Compared to the Unruh channel, we have now constrained Alice’s encoding operations. The receiver Bob still sees the right Rindler wedge defined by the half-space $B$.
@@
$\max_{\psi\in\mathcal{H}_{A}\otimes\mathcal{H}_{R}}I(B:R)_{\mathcal{M}(\psi)}\leq 2\pi L\langle H\rangle_{\mathcal{M}(\psi^{*})}=:2\pi LE\ ,$ (5.4)
@@
Since we have considered the case of restraining only the decoder and the case of restraining both the encoder and the decoder, one could ask what if only the encoder is restrained. It turns out that it is still crucial to restrain the decoder for the Bekenstein bound to hold. We revisit an argument due to Page in Appendix B, which essentially shows that no Bekenstein bound could work otherwise.
@@ Appendix A No Bekenstein bound for $\alpha$-bits
Hayden-Penington proposed a generalization of a zero-bit, called an $\alpha$-bit ($0\leq\alpha\leq 1$hayden2020approximate . An $\alpha$-bit interpolates between a zero-bit ($\alpha=0$) and a qubit ($\alpha=1$). The intuition is that while a noisy channel might only be able to transmit a decent amount of noisy qubits, we can still make good use of them by distilling an $\alpha$-fraction of each qubit. Just like the zero-bits, the so-obtained $\alpha$-bits can be used as tighter communications resources in various tasks. [^28] For instance, the $\alpha$-bit dense-coding resource inequality reads, $1\;\alpha$-bit $+1\;\mathrm{ebit}\geq(1+\alpha)\;\mathrm{cbits}$.
Let us define $\alpha$-bits in the communication scenario. (cf. Section 3 of hayden2020approximate for the precise definition.) Recall that the zero-bit capacity $Q_{0}(\mathcal{M})$ characterizes how much information the channel can transmit if we only require Bob to decode a two-dimensional (or any constant) subspace. The $\alpha$-bit capacity $Q_{\alpha}(\mathcal{M})$ is defined similarly by requiring Bob to decode, with sufficiently small error at large enough $n$, a $k^{n\alpha}$-dimensional subspace of the $k^{n}$-dimensional input message transmitted by $\mathcal{M}^{\otimes n}$. We then say that a channel $\mathcal{M}$ transmits a $Q_{\alpha}(\mathcal{M})$ amount of $\alpha$-bits (asymptotically).
@@
It is natural to ask for which $\alpha$ values there is a violation of the Bekenstein bound $\langle K_{B}\rangle=\beta E$. Given a fixed $\beta\omega$, we find that there is a violation for
@@
As a sanity check, we note that there is no lower bound for the range of Bekenstein-violating $\alpha$ values, so the zero-bits always violate the bound at some temperature. On the other hand, the upper bounds are strictly less than $1$ because $Q(\mathcal{N})<\beta E$, meaning that qubits obey the Bekenstein bound. Also, the maximal $\alpha$ that satisfies (A.3) can be arbitrarily close to $1$ at a sufficiently large $d$, because in this limit $Q(\mathcal{N})$ tends to $\beta E$ from below. In Fig. 5, we plotted a “phase diagram” that illustrates (A.3) for $d=2^{6}$.
@@ Appendix D Zero-bit dense coding
$X:=\sum_{0\leq j<d}\left|j+1\ \text{mod}\ d\right\rangle\!\!\left\langle j\right|,\quad Z:=\sum_{0\leq j<d}e^{i2\pi j/d}\left|j\right\rangle\!\!\left\langle j\right|$ (D.1)